OpenStudy (anonymous):

If a>b and a(b-a)=0, which one or more of the following must be true? i) a=0 ii) b<0 iii) a-b>0

hartnn (hartnn):

if a(b-a) = 0 then either a=0 OR b-a =0 did u get this?

hartnn (hartnn):

@Ventricate did u get how i got that?

hartnn (hartnn):

anyways , lets continue b-a=0 gives b=a but its given that a>b so b-a=0 is NOT true hence a=0 must be true

OpenStudy (anonymous):

ya i get that but the answer key says all three is true

hartnn (hartnn):

since a>b and a=0 then b must be negative so b<0

hartnn (hartnn):

-b>0 (u flip the sign when you multiply by a negative quantity) add a to both sides a-b>a and since a=0 a-b > 0

OpenStudy (anonymous):

ahh makes more sense thank you!

hartnn (hartnn):

welcome ^_^